OpenStudy (anonymous):

Solve the system by substitution 2x-y+z=-4 z=5 -2x+3y-z=-10

OpenStudy (solomonzelman):

Add `2x-y+z=-4` and `-2x+3y-z=-10` to each other. This will eliminate x and z variables.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2x{~~~~}-y{~~~~}+z{~~~}={~~}-4 }\) \(\Large\color{red}{+}\) \(\large\color{black}{ \displaystyle -2x{~~}+3y{~~}-z{~~~}={~~}-10 }\) \(\large\color{blue }{ \displaystyle ^\text{_________________________________} }\)

OpenStudy (anonymous):

Well I have multiple choice A.(-8,7,5) B.(-8,-7,5) C.(8,-7,5) D.(-8,-7,-5) I think it is C?... What do you think? @SolomonZelman

OpenStudy (solomonzelman):

I try not to think about the options, rather help you solve the equation.

OpenStudy (anonymous):

okay could you please hold on a sec while I solve it @SolomonZelman

OpenStudy (anonymous):

I got... x+2y=-14 @SolomonZelman

OpenStudy (solomonzelman):

2x+(-2x)= what?

OpenStudy (solomonzelman):

4 - 4 = 0 (Quantity) - (Quantity) = 0

OpenStudy (anonymous):

2x+(-2)=0x

OpenStudy (solomonzelman):

yes, so you should get: 2y=-14, from the addition of the two equations

OpenStudy (solomonzelman):

Can you solve for y?

OpenStudy (anonymous):

okay so then would I divide 2 from -14 to get y? -14/2=-7

OpenStudy (solomonzelman):

yes, so y=-7, and you are given that z=5.

OpenStudy (solomonzelman):

use the first equation: 2x-y+z=-4 and use the fact that you know that y=-7 and z=5, to solve for x.

OpenStudy (anonymous):

I got x=-8

OpenStudy (anonymous):

x=-8,y=-7, and z=5

OpenStudy (solomonzelman):

2x-(-7)+5=-4 2x+7+5=-4 2x+12=-4 2x=-16 x=-8 Yes, beautiful !!

OpenStudy (solomonzelman):

And that is indeed the correct final answer:

OpenStudy (anonymous):

Yay! Thank you for your help :)